OpenStudy (anonymous):

how many cubic yards are there in 13,365 ft cubed ?

OpenStudy (help!!!!):

convert yards to feet

OpenStudy (anonymous):

how do I do that?

OpenStudy (tkhunny):

Multiply by 1 (one)!! \(13365\;ft^{3}\cdot \left(\dfrac{1\;yd}{3\;ft}\right)^{3}\)

OpenStudy (anonymous):

the answer is one?

OpenStudy (anonymous):

13,365/27 = 495 yd^3

OpenStudy (anonymous):

thanks so much!(:

OpenStudy (tkhunny):

No, the answer it not 1. Multiply by 1, as demonstrated.